OpenStudy (anonymous):

write an equation in point-slope form of the line through point j(-5,6) with slope -4

OpenStudy (anonymous):

the form is simply \[(y-y_1)=m(x-x_1)\]just plug-in the point(-5,6) and slope m=-4

OpenStudy (anonymous):

(y-y1)=m(x-x1) You now have to plug in your values to the equation for point-slope form. (y-6)=-4(x+5)

OpenStudy (anonymous):

What do I do after that? Sorry im very slow when it comes to math.

OpenStudy (anonymous):

That's it! (y-6)=-4(x+5) is the equation in point-slope form!

OpenStudy (anonymous):

Thank you so much! I have a test tomorrow and that stuff is on there for "review"